# Continuous operators from spaces of Lipschitz functions

### v1

@@ 5. Operators from spaces $\operatorname{Lip}_{0}(M)$ onto $\ell_{1}$
In this section we will study the question for which metric spaces $M$ there exists a continuous linear surjection from the space $\operatorname{Lip}_{0}(M)$ onto the Banach spaces $c_{0}$ or $\ell_{1}$. Recall that Rosenthal [58, Remark 2] proved that the Banach space $\ell_{\infty}(\kappa)$, where $\kappa$ is an infinite cardinal number, always admits a continuous linear surjection onto the Hilbert space $\ell_{2}(2^{\kappa})$, so in particular onto the space $\ell_{2}$. Since for any infinite metric space $M$, by Theorem 3.2, $\operatorname{Lip}_{0}(M)$ contains a complemented copy of $\ell_{\infty}(d(M))$, Rosenthal’s result yields the following observation, being a main motivation for this section.
@@
Note that there is no continuous linear surjection from $\ell_{2}$ onto $c_{0}$ or any space $\ell_{p}$ with $p\neq 2$ (since they are not Hilbert spaces). Note also that for any infinite metric space $M$ the Lipschitz-free space $\mathcal{F}(M)$ and the dual space $\operatorname{Lip}_{0}(M)^{*}$ both contain complemented copies of $\ell_{1}$ (in the latter case by the fact that the dual space $\ell_{\infty}^{*}$ contains such a copy). Since for any Banach space $X$ the dual space $X^{*}$ is complemented in the third dual space $X^{***}$ and so in the $(2n+1)$-th dual space $X^{*(2n+1)}$ for any $n\geq 0$, it follows that, again for any $n\geq 0$, the $n$-th dual space $\operatorname{Lip}_{0}(M)^{*(n)}$ contains a complemented copy of $\ell_{1}$ if $n$ is odd, and a complemented copy of $\ell_{\infty}$ if $n$ is even.
@@
Proof.
For any Banach space $X$, by [49, Theorem 2] or [64, Corollary 5.4], the space $\operatorname{Lip}_{0}(X)$ contains a $1$-complemented copy of $X^{*}$. Thus, by the discussion preceding the proposition, for all $n\geq 0$, $\operatorname{Lip}_{0}(E^{*(2n)})$ admits a continuous operator onto $\ell_{1}$. ∎
@@
If $E$ is a Banach space with no non-trivial cotype, then $\operatorname{Lip}_{0}(E)$ admits a continuous linear surjection onto $\ell_{1}$.
@@
Since separable Banach spaces have the lifting property, see [29, Theorem 3.1], $E$ and hence also $F$ embeds linearly into $\mathcal{F}(E)$, by [29, Theorem 2.12]. By transposition, we obtain a continuous linear surjection $\operatorname{Lip}_{0}(E)\rightarrow\ell_{1}$. The second statement follows from [29, Corollary 3.3] saying that a Banach space containing an isometric copy of $c_{0}$ also contains a linear subspace isometrically isomorphic to $c_{0}$. ∎
Recall that a metric space $M$ is an absolute Lipschitz retract if $M$ is a Lipschitz retract of every metric space containing $M$. Equivalently, a metric space $M$ is an absolute Lipschitz retract if for all metric spaces $P\subseteq N$ and every Lipschitz mapping $f\colon P\rightarrow M$ there is a Lipschitz extension $F\colon N\rightarrow M$ of $f$. This characterization immediately implies that bilipschitz copies of absolute Lipschitz retracts are absolute Lipschitz retracts. Lindenstrauss [49] proved that Banach spaces $c_{0}$ and $C(K)$ for $K$ metric compact are absolute Lipschitz retracts. We refer the reader to [10, Section 1.1] for an introduction to absolute Lipschitz retracts.
@@
Since $c_{0}$ is separably injective, $E$ is isomorphic to $E\oplus\mathbb{R}$, so by the main result of  [15] the spaces $\operatorname{Lip}_{0}(S_{E})$ and $\operatorname{Lip}_{0}(E)$ are isomorphic. Since $E$ contains an isomorphic copy of $c_{0}$, Proposition 5.4 yields that there is a continuous linear surjection from $\operatorname{Lip}_{0}(E)$ onto $\ell_{1}$ and hence from $\operatorname{Lip}_{0}(S_{E})$ onto $\ell_{1}$.
@@
Note that in Theorem 5.6 we may require that $E$ contains an isomorphic copy of a predual of $\ell_{1}$, as every such predual contains a copy of $c_{0}$ (by Johnson and Zippin [40]). Note also that Benyamini and Lindenstrauss [9] constructed a predual of $\ell_{1}$ which is not isomorphic to a complemented subspace of any space $C(K)$ (though by [40, Theorem] it is isomorphic to a quotient of the space $C(2^{\mathbb{N}})$, where $2^{\mathbb{N}}$ denotes the Cantor space).
Proposition 5.9.
@@
By the result of Dalet [23, Proposition 79], the space $\operatorname{Lip}_{0}(\ell_{1})$ contains a complemented copy of $\ell_{1}$. Since a continuous linear projection is in particular a Lipschitz retraction, the conclusion follows from Proposition 5.5. ∎
Using the above results, we now provide a list of several examples of Banach spaces $E$ such that $\operatorname{Lip}_{0}(E)$ has a quotient isomorphic to $\ell_{1}$. As usual, by $\mathcal{B}(H)$ we denote the C*-algebra of all bounded operators on a Hilbert space $H$, by $\mathcal{L}(Y,Z)$ the Banach space of all continuous operators from a normed space $Y$ into a Banach space $Z$, and by $\mbox{Bil}(Y\times W,Z)$ the Banach space of all continuous bilinear mappings from the product $Y\times W$ of Banach spaces $Y$ and $W$ into a Banach space $Z$.
@@
    $E=L_{\infty}(\Omega,\Sigma,\mu)^{*(n)}$ for some measure space $(\Omega,\Sigma,\mu)$,
@@
    $E=\mathcal{B}(H)^{*(n)}$ for some Hilbert space $H$,
@@
    $E$ is a predual of a C*-algebra,
  6. (6)
    $E=L_{1}(\Omega,\Sigma,\mu)$ for some measure space $(\Omega,\Sigma,\mu)$,
@@
(1) If $E$ is an infinite-dimensional C*-algebra, then it is well-known that $E$ contains an infinite-dimensional maximal abelian subalgebra $M$. By the classical Gelfand–Naimark theorem $M$ is isomorphic to the space $C_{0}(X)$ for some infinite locally compact space $X$. Consequently, $M$ contains a copy of $c_{0}$ and so $E$ contains such a copy. By Corollary 5.7 we get a continuous operator from $\operatorname{Lip}_{0}(E)$ onto $\ell_{1}$.
Assume that $E=X^{*(n)}$ for some infinite-dimensional C*-algebra $X$ and $n>0$. If $n$ is even, then $E$ carries a compatible C*-algebraic structure and so, by the previous argument, $E$ contains a copy of $c_{0}$. If $n$ is odd, then $c_{0}$ can be embedded into $X^{*(n+1)}=E^{*}$ and hence, by [59, Corollary 1.2], $E$ contains a complemented copy of $\ell_{1}$. In either case, referring to Corollary 5.7 or to Proposition 5.9, we get a continuous operator from $\operatorname{Lip}_{0}(E)$ onto $\ell_{1}$.
(2), (3), and (4) follow from (1).
(5) By an argument as in the second part of (1), we get that $E$ contains a complemented copy of $\ell_{1}$ and we use Proposition 5.9.
(6) follows from (5).
(7) By [25, Theorem 1], $E$ contains a copy of $c_{0}$, so we use Corollary 5.7.
@@
(9) If $E=\operatorname{Lip}_{0}(M)^{*(n)}$ and $n$ is even, then, by Theorem 3.2 and the discussion after Corollary 5.1, $E$ contains a copy of $\ell_{\infty}$ and so, in particular, a copy of $c_{0}$; we then use Corollary 5.7. If $n$ is odd, then, similarly, $E$ contains a complemented copy of $\ell_{1}$, so we use Proposition 5.9.
(10) follows from Corollary 3.3 and Proposition 5.9. ∎
@@
Recall that for a metric space $M$ a function $f\colon M\rightarrow\mathbb{R}$ is called locally flat if for every $x\in M$ and every $\varepsilon>0$ there is $\delta>0$ such that $|f(y)-f(z)|\leq\varepsilon\rho(y,z)$ for all points $y,z\in B(x,\delta)$. For a compact metric space $M$ with the base point, the space $\mathrm{lip}_{0}(M)$ is defined as the subspace of $\operatorname{Lip}_{0}(M)$ consisting of all locally flat functions. For non-compact metric spaces $M$ additional flatness assumptions are required in the definition of $\mathrm{lip}_{0}(M)$, see [64, Definition 4.15]. The space $\mathrm{lip}_{0}(M)$ is said to separate points uniformly if there is a constant $a\in(0,1]$ such that for all $x,y\in M$ there is a function $f\in\mathrm{lip}_{0}(M)$ with $\|f\|_{\operatorname{Lip}_{0}(M)}\leq 1$ and $|f(x)-f(y)|=a\rho(x,y)$. The class of all metric spaces $M$ for which the spaces $\mathrm{lip}_{0}(M)$ separate points uniformly contains e.g. the Cantor space $2^{\mathbb{N}}$ and all countable compact spaces, see [64, Section 4.1].
@@
By [64, Theorem 4.38], if a metric space $M$ is infinite and the space $\mathrm{lip}_{0}(M)$ separates points uniformly, then the dual space of $\mathrm{lip}_{0}(M)$ is isometrically isomorphic to $\mathcal{F}(M)$, and so, since $\mathcal{F}(M)$ contains a complemented copy of $\ell_{1}$, by Proposition 5.2, $\operatorname{Lip}_{0}(\mathrm{lip}_{0}(M))$ admits a continuous operator onto $\ell_{1}$.
@@
There is an infinite-dimensional arcwise-connected $\sigma$-compact complete metric space $M$ which does not contain any bilipschitz copy of the unit sphere of an infinite-dimensional normed space but $\operatorname{Lip}_{0}(M)$ still admits a continuous operator onto $\ell_{1}$.
@@
We consider the metric space $M=\bigsqcup_{n\in\mathbb{N}}\ell_{\infty}^{n}$ which is defined in accordance with [64, Definition 1.13] as follows: it is the quotient space of the disjoint union $\bigsqcup_{n\in\mathbb{N}}\ell_{\infty}^{n}$ with all the base points of the spaces $\ell_{\infty}^{n}$ identified as the new single base point, and for points $x\in\ell_{\infty}^{n}$ and $y\in\ell_{\infty}^{m}$ their distance is defined as
@@
Note that $M$ is an arcwise-connected $\sigma$-compact complete metric space which does not contain any homeomorphic copy of the unit sphere of an infinite-dimensional normed space. By [64, Proposition 2.8], we have the isomorphism

### v3

@@ 5. Operators from spaces $\operatorname{Lip}_{0}(M)$ onto $\ell_{1}$
In this section we will study the question for which metric spaces $M$ there exists a continuous linear surjection from the space $\operatorname{Lip}_{0}(M)$ onto the Banach spaces $c_{0}$ or $\ell_{1}$. Recall that Rosenthal [36, Theorem 4.22] proved that the Banach space $\ell_{\infty}(\kappa)$, where $\kappa$ is an infinite cardinal number, always admits a continuous linear surjection onto the Hilbert space $\ell_{2}(2^{\kappa})$, so in particular onto the space $\ell_{2}$. Since for any infinite metric space $M$, by Theorem 3.2, $\operatorname{Lip}_{0}(M)$ contains a complemented copy of $\ell_{\infty}(d(M))$, Rosenthal’s result yields the following observation, being a main motivation for this section.
@@
Note that there is no continuous linear surjection from $\ell_{2}$ onto $c_{0}$ or any space $\ell_{p}$ with $p\neq 2$ (since they are not Hilbert spaces). Note also that for any infinite metric space $M$ the Lipschitz-free space $\mathcal{F}(M)$ and the dual space $\operatorname{Lip}_{0}(M)^{*}$ both contain complemented copies of $\ell_{1}$ (in the latter case by Theorem 3.2 and the fact that the dual space $\ell_{\infty}^{*}$ contains such a copy). It follows that, for any $n\geq 0$, the $n$-th dual space $\operatorname{Lip}_{0}(M)^{*(n)}$ contains a complemented copy of $\ell_{1}$ if $n$ is odd, and a complemented copy of $\ell_{\infty}$ if $n$ is even.
Recall that, for any Banach space $E$, the dual space $E^{*}$ is complemented in the third dual space $E^{***}$ and so in the $(2n+1)$-th dual space $E^{*(2n+1)}$ for any $n\geq 0$, and that, by [52, Theorem 2] or [66, Corollary 5.4], the space $\operatorname{Lip}_{0}(E)$ contains a $1$-complemented copy of $E^{*}$. We thus have the following result.
@@
If $E$ is a separable Banach space with no non-trivial cotype, then $\operatorname{Lip}_{0}(E)$ admits a continuous linear surjection onto $\ell_{1}$.
@@
Since separable Banach spaces have the lifting property, see [31, Theorem 3.1], $E$ and hence also $F$ embeds linearly into $\mathcal{F}(E)$, by [31, Theorem 2.12]. By transposition, we obtain a continuous linear surjection $\operatorname{Lip}_{0}(E)\rightarrow\ell_{1}$. The second statement follows from [31, Corollary 3.3] saying that a Banach space containing an isometric copy of $c_{0}$ also contains a linear subspace isometrically isomorphic to $c_{0}$. ∎
Recall that a metric space $M$ is an absolute Lipschitz retract if $M$ is a Lipschitz retract of every metric space containing $M$. Equivalently, a metric space $M$ is an absolute Lipschitz retract if for all metric spaces $P\subseteq N$ and every Lipschitz mapping $f\colon P\rightarrow M$ there is a Lipschitz extension $F\colon N\rightarrow M$ of $f$. This characterization immediately implies that bilipschitz copies of absolute Lipschitz retracts are absolute Lipschitz retracts. Lindenstrauss [52] proved that Banach spaces $c_{0}$ and $C(K)$ for $K$ metric compact are absolute Lipschitz retracts. We refer the reader to [9, Section 1.1] for an introduction to absolute Lipschitz retracts.
@@
Since $c_{0}$ is separably injective, $E$ is isomorphic to $E\oplus\mathbb{R}$, so by the main result of  [16] the spaces $\operatorname{Lip}_{0}(S_{E})$ and $\operatorname{Lip}_{0}(E)$ are isomorphic. Since $E$ contains an isomorphic copy of $c_{0}$, Lemma 5.4 yields that there is a continuous linear surjection from $\operatorname{Lip}_{0}(E)$ onto $\ell_{1}$ and hence from $\operatorname{Lip}_{0}(S_{E})$ onto $\ell_{1}$.
@@
Note that in Theorem 5.6 we may require that $E$ contains an isomorphic copy of a predual of $\ell_{1}$, as every such predual contains a copy of $c_{0}$ (by Johnson and Zippin [42]). Note also that Benyamini and Lindenstrauss [8] constructed a predual of $\ell_{1}$ which is not isomorphic to a complemented subspace of any space $C(K)$ (though by [42, Theorem] it is isomorphic to a quotient of the space $C(2^{\mathbb{N}})$, where $2^{\mathbb{N}}$ denotes the Cantor space).
Corollary 5.9.
@@
By the result of Dalet [24, Proposition 79], the space $\operatorname{Lip}_{0}(\ell_{1})$ contains a complemented copy of $\ell_{1}$. Since a continuous linear projection is in particular a Lipschitz retraction, the conclusion follows from Proposition 5.5. ∎
Using the above results, we now provide a list of several examples of Banach spaces $E$ such that $\operatorname{Lip}_{0}(E)$ has a quotient isomorphic to $\ell_{1}$. Below by $\mathcal{L}(Y,Z)$ we denote the Banach space of all continuous operators from a normed space $Y$ into a Banach space $Z$, and by $\mbox{Bil}(Y\times W,Z)$ the Banach space of all continuous bilinear mappings from the product $Y\times W$ of Banach spaces $Y$ and $W$ into a Banach space $Z$.
@@
    $E=L_{\infty}(\Omega,\Sigma,\mu)^{*(n)}$ for some $\sigma$-finite measure space $(\Omega,\Sigma,\mu)$,
@@
    $E$ is a predual space of the space $C(K)$ for some compact space $K$,
@@
    $E=L_{1}(\Omega,\Sigma,\mu)$ for some $\sigma$-finite measure space $(\Omega,\Sigma,\mu)$,
@@
(1) Recall that the space $C_{0}(Y)$ contains a copy of $c_{0}$, the dual space $C_{0}(Y)^{*}$ contains a complemented copy of $\ell_{1}$ and hence that the bidual space $C_{0}(Y)^{**}$ contains a complemented copy of $\ell_{\infty}$. It follows that, for each odd $m\in\mathbb{N}$, the space $C_{0}(Y)^{*(m)}$ contains a complemented copy of $\ell_{1}$ and the space $C_{0}(Y)^{*(m+1)}$ contains a complemented copy of $\ell_{\infty}$ and so a copy of $c_{0}$. Consequently, if $n$ is even, we use Corollary 5.7, and if $n$ is odd, we appeal to Corollary 5.9.
(2) follows from [1, Example 4.2.9] and (1).
(3) Since $C(K)$ contains a copy of $c_{0}$ and is a dual space, by [60, Corollary 1.2] $E$ contains a complemented copy of $\ell_{1}$ and so we use Corollary 5.9.
(4) follows from [1, Example 4.2.9] and (3).
(5) By [27, Theorem 1], $E$ contains a copy of $c_{0}$, so we use Corollary 5.7.
@@
(7) If $E=\operatorname{Lip}_{0}(M)^{*(n)}$ and $n$ is even, then, by the discussion after Corollary 5.1, $E$ contains a copy of $\ell_{\infty}$ and so, in particular, a copy of $c_{0}$; we then use Corollary 5.7. If $n$ is odd, then, similarly, $E$ contains a complemented copy of $\ell_{1}$, so we use Corollary 5.9.
(8) follows from Corollaries 3.3 and 5.9. ∎
Even though in this paper we work exclusively with real Banach spaces, the arguments presented in the proof of Theorem 5.10 can be generalized to the class of complex C*-algebras. However, while in the below proof we use the fact that the C*-algebras are complex to obtain either $c_{0}$ or $\ell_{1}$ as a (complemented) subspace of $E$, for the space $\operatorname{Lip}_{0}(E)$ (consisting of real-valued functions), we consider $E$ as a real Banach space by forgetting its additional complex structure. In fact, in the cases where we exhibit a copy of $c_{0}$ in $E$, taking into account Corollary 5.7, we use only that $E$ is a metric space containing a bilipschitz copy (of the unit sphere) of $c_{0}$. Note that the following result in particular applies to the C*-algebras $X=\mathcal{B}(H)$ of all bounded operators on infinite-dimensional Hilbert spaces $H$.
Theorem 5.11.
Let $n\geq 0$. Let $E$ be an infinite-dimensional complex Banach space. If $E=X^{*(n)}$ for some C*-algebra $X$ or $E$ is a predual of a C*-algebra, then $\operatorname{Lip}_{0}(E)$ admits a continuous operator onto $\ell_{1}$.
Proof.
Assume first that $E$ is an infinite-dimensional C*-algebra. Then, it is well-known that $E$ contains an infinite-dimensional maximal abelian subalgebra $M$ (see [43, Exercise 4.6.12]). By the classical Gelfand–Naimark theorem $M$ is isomorphic to the complex space $C_{0}(X)$ for some infinite locally compact space $X$. Consequently, $M$ contains a copy of the real Banach space $c_{0}$ and so $E$ contains such a copy, too. By Corollary 5.7 we get a continuous operator from $\operatorname{Lip}_{0}(E)$ onto $\ell_{1}$.
Assume that $E=X^{*(n)}$ for some infinite-dimensional C*-algebra $X$ and $n>0$. If $n$ is even, then $E$ carries a compatible C*-algebraic structure (see [61, Section 1.17]) and so, by the previous argument, $E$ contains a copy of the real Banach space $c_{0}$. If $n$ is odd, then similarly the complex Banach space $c_{0}$ can be embedded into complex $X^{*(n+1)}=E^{*}$ and hence, by [60, Corollary 1.2], the complex Banach space $E$ contains a complemented copy of the complex Banach space $\ell_{1}$. Consequently, the real Banach space $E$ contains a complemented copy of the real Banach space $\ell_{1}$. In either case, referring to Corollary 5.7 or to Corollary 5.9, we get a continuous operator from $\operatorname{Lip}_{0}(E)$ onto $\ell_{1}$.
If $E$ is a predual of a C*-algebra, then we proceed similarly as in the previous paragraph for odd $n$ to show that $E$ contains a complemented copy of $\ell_{1}$ and we use Corollary 5.9. ∎
@@
Recall that for a metric space $M$ a function $f\colon M\rightarrow\mathbb{R}$ is called locally flat if for every $x\in M$ and every $\varepsilon>0$ there is $\delta>0$ such that $|f(y)-f(z)|\leq\varepsilon\rho(y,z)$ for all points $y,z\in B(x,\delta)$. For a compact metric space $M$ with the base point, the space $\mathrm{lip}_{0}(M)$ is defined as the subspace of $\operatorname{Lip}_{0}(M)$ consisting of all locally flat functions. For non-compact metric spaces $M$ additional flatness assumptions are required in the definition of $\mathrm{lip}_{0}(M)$, see [66, Definition 4.15]. The space $\mathrm{lip}_{0}(M)$ is said to separate points uniformly if there is a constant $a\in(0,1]$ such that for all $x,y\in M$ there is a function $f\in\mathrm{lip}_{0}(M)$ with $\|f\|_{\operatorname{Lip}_{0}(M)}\leq 1$ and $|f(x)-f(y)|=a\rho(x,y)$. The class of all metric spaces $M$ for which the spaces $\mathrm{lip}_{0}(M)$ separate points uniformly contains e.g. the Cantor space $2^{\mathbb{N}}$ and all countable compact spaces, see [66, Section 4.1].
@@
By [66, Theorem 4.38], if a metric space $M$ is infinite and the space $\mathrm{lip}_{0}(M)$ separates points uniformly, then the dual space of $\mathrm{lip}_{0}(M)$ is isometrically isomorphic to $\mathcal{F}(M)$, and so, since $\mathcal{F}(M)$ contains a complemented copy of $\ell_{1}$, by Proposition 5.2, $\operatorname{Lip}_{0}(\mathrm{lip}_{0}(M))$ admits a continuous operator onto $\ell_{1}$.
@@
There is an arcwise-connected $\sigma$-compact complete metric space $M$ of infinite covering dimension which does not contain any bilipschitz copy of the unit sphere of an infinite-dimensional normed space but $\operatorname{Lip}_{0}(M)$ still admits a continuous operator onto $\ell_{1}$.
@@
We consider the metric space $M=\bigsqcup_{n\in\mathbb{N}}\ell_{\infty}^{n}$ which is defined in accordance with [66, Definition 1.13] as follows: it is the quotient space of the disjoint union $\bigsqcup_{n\in\mathbb{N}}\ell_{\infty}^{n}$ with all the base points of the spaces $\ell_{\infty}^{n}$ identified as the new single base point, and for points $x\in\ell_{\infty}^{n}$ and $y\in\ell_{\infty}^{m}$ their distance is defined as
@@
Note that $M$ is an arcwise-connected $\sigma$-compact complete metric space. Since for each $n\in\mathbb{N}$ the space $M$ contains $\ell_{\infty}^{n}$ as a closed subset, it has infinite covering dimension, see e.g. [26, Theorem 3.1.4].
We claim that $M$ contains no homeomorphic copy of the unit sphere of an infinite-dimensional normed space. To see this, note first that the sphere $S$ of an infinite-dimensional normed space has infinite covering dimension and does not contain any point $x$ such that $S\setminus\{x\}$ is disconnected, so every homeomorphic copy of $S$ would have to be contained in one of the $n$-dimensional pieces $\ell_{\infty}^{n}$ of $M$. Since homeomorphisms preserve the covering dimension, see e.g. [26, p. 54], this would result in a closed infinite-dimensional subset of $\ell_{\infty}^{n}$ for some $n$, which is impossible.
By [66, Proposition 2.8], we have the isomorphism